Calculate the number of moles of solute needed to prepare 1800 mL of 1.2 mol/L HCl, hydrochloric acid solution. 


1
Expert's answer
2021-03-25T04:09:19-0400

Molar Mass of HCL = 36.458 g/mol 1.81 of 1.2mol/L = 1.2mol/L HCl/1.8 L = 0.66667 mol

1 M = 36.458 g

0.6667M =?

= 36.458× 0.6667 = 24.3053g

= 1800/24.3053 = 74.0579 moles